### v1

@@ 1.2.4 The formulation of generic projection and unconstrained PPE (GePUP)
Inspired by Liu, Liu & Pego [20], Zhang [5] proposed the GePUP formulation of the INSE based on a commutator of the generic projection and the Laplacian. In order to accommodate the fact that the discrete velocity might not be divergence-free, the evolutionary variable of the GePUP formulation (21) is designed to be a non-solenoidal velocity $\mathbf{w}={\mathcal{P}}\mathbf{u}$ instead of the divergence-free velocity $\mathbf{u}$ in the UPPE formulation (19). More precisely, the evolutionary variable $\mathbf{u}$ in the time-derivative term $\frac{\partial\mathbf{u}}{\partial t}$ is perturbed to a non-solenoidal velocity $\mathbf{w}:=\mathbf{u}-\nabla\phi$ where $\phi$ is some scalar function; meanwhile in the diffusion term we change $\mathbf{u}$ to $\mathbf{w}$ to set up a mechanism that drives the divergence towards zero. Then, there is no need to worry about the influence of $\nabla\cdot\mathbf{w}\neq 0$ on numerical stability because the evolution of $\mathbf{w}$ is not subject to the divergence-free constraint. These ideas lead to the GePUP formulation:
$\frac{\partial\mathbf{w}}{\partial t}$ $=\mathbf{g}-\mathbf{u}\cdot\nabla\mathbf{u}-\nabla q+\nu\Delta\mathbf{w}$ in $\text{in }\Omega,$ (21a)
$\mathbf{w}$ $=\mathbf{0},\ \ \mathbf{u}\cdot\bm{\tau}=0,$ on $\text{on }\partial\Omega,$ (21b)
$\mathbf{u}$ $={\mathscr{P}}\mathbf{w}$ in $\text{in }\Omega,$ (21c)
$\mathbf{u}\cdot\mathbf{n}$ $=0$ on $\text{on }\partial\Omega,$ (21d)
$\Delta q$ $=\nabla\cdot(\mathbf{g}-\mathbf{u}\cdot\nabla\mathbf{u})$ in $\text{in }\Omega,$ (21e)
$\mathbf{n}\cdot\nabla q$ $=\mathbf{n}\cdot(\mathbf{g}+\nu\Delta\mathbf{u}-\nu\nabla\nabla\cdot\mathbf{u})$ on $\text{on }\partial\Omega.$ (21f)
By (21), $\mathbf{w}$ satisfies the heat equation
$\frac{\partial\left(\nabla\cdot\mathbf{w}\right)}{\partial t}=\nu\Delta\left(\nabla\cdot\mathbf{w}\right)\qquad\text{ in }\Omega,$ (22)
which, combined with $\nabla\cdot\mathbf{w}=0$ or $\mathbf{n}\cdot\nabla\nabla\cdot\mathbf{w}=0$ on $\partial\Omega$, drives $\nabla\cdot\mathbf{w}$ towards zero.
The GePUP formulation (21) not only retains the advantages of UPPE, but also provides certain degrees of freedom in the design of high-order projection methods.
@@ 1.3 The contribution of this work
In this work, we couple GePUP with the electric boundary conditions [22, 23] and the scalar auxiliary variable (SAV) approach [24, 25] to propose GePUP-E and GePUP-ES, variants of GePUP that enforce the solenoidal conditions, preserve energy stability, decouple time integration from spatial discretization, and lead to versatile algorithms that are fourth-order accurate both in time and in space. The letter ‘E’ in the acronyms stands for the electric boundary conditions while the letter ‘S’ for the SAV approach; altogether ‘ES’ also stands for energy stability.
GePUP-E and GePUP-ES answer all the challenges (A–D) posed in Section 1.1.
  1. [(CTB-1)]
  2. 1.
    We establish the GePUP-E formulation in (39) on solid ground by proving their equivalence to the INSE, the convergence of the non-solenoidal velocity $\mathbf{w}$ to the divergence-free velocity $\mathbf{u}$, the exponential decay of the divergence $\nabla\cdot\mathbf{w}$, and the monotonic decrease of the kinetic energy. GePUP-E resolves the difficulties in (RSN-1,2).
  3. 2.
    By coupling GePUP-E to SAV [24, 25], we propose the GePUP-ES formulation in (54), design a family of semi-discrete GePUP-ES algorithms, and prove their energy stability in Theorem 10.
  4. 3.
    Based on (CTB-2), we further propose a family of fully discrete INSE solvers, named GePUP-ES-SDIRK, to answer all challenges in (A–D).
(CTB-1,2,3) are explained in Sections 3, 4, and 5, respectively. In Section 2, we introduce notation to make this paper somewhat self-contained. We test GePUP-ES-SDIRK in Section 6 and draw conclusions in Section 7.
@@ 2 Preliminaries
[…]
2.1 The Leray-Helmholtz projection ${\mathscr{P}}$
This subsection hinges on the following well-known result on boundary value problems (BVPs).
Theorem 1 (Solvability of BVPs with pure Neumann conditions).
@@ 2.1 The Leray-Helmholtz projection ${\mathscr{P}}$
Proof.
See [26, page 409]. ∎
Theorem 2 (Helmholtz decomposition).
@@
since Theorem 1 uniquely determines $\nabla\phi$ with $\oint_{\partial\Omega}\mathbf{v}\cdot\mathbf{n}=0$. ∎
@@
The proof of Theorem 2 implies the constructive form
${\mathscr{P}}={\mathcal{I}}-\nabla(\Delta_{n})^{-1}\nabla\cdot,$
where $(\Delta_{n})^{-1}$ denotes solving (25). For a ${\mathcal{C}}^{1}$ vector field $\mathbf{v}^{*}$ and a ${\mathcal{C}}^{1}$ scalar field $\phi$, we have
@@ 2.2 The Laplace-Leray commutator $\Delta{\mathscr{P}}-{\mathscr{P}}\Delta$
On periodic domains, $\Delta$ and ${\mathscr{P}}$ commute. However, one main difficulty for no-slip domains is the fact that $\Delta{\mathscr{P}}-{\mathscr{P}}\Delta\neq\mathbf{0}$. In this subsection we rephrase several results in [20].
Lemma 1.
@@
satisfies $\nabla\cdot{\mathcal{B}}=0$, $\Delta{\mathscr{P}}={\mathcal{B}}$, and in three dimensions ${\mathcal{B}}=-\nabla\times\nabla\times$.
Proof.
@@
Then Lemma 1 and (30) give
Corollary 1.
@@
$[\Delta,{\mathscr{P}}]:=\Delta{\mathscr{P}}-{\mathscr{P}}\Delta=({\mathcal{I}}-{\mathscr{P}})\Delta{\mathscr{P}}=({\mathcal{I}}-{\mathscr{P}}){\mathcal{B}}=-({\mathcal{I}}-{\mathscr{P}})(\nabla\times\nabla\times),$ (31)
@@ 2.3 The Stokes pressure
By (31), the action of the Laplace-Leray commutator on any vector field $\mathbf{v}^{*}$ results in the gradient of some scalar field. In the case of $\mathbf{v}^{*}$ being the velocity $\mathbf{u}$ in the INSE, the scalar is known as the Stokes pressure [20]:
@@
Then (27) and $\nabla\cdot{\mathcal{B}}=0$ in Lemma 1 yield $\Delta p_{s}=0$, i.e., the Stokes pressure is harmonic. Interestingly, the vector field $\nabla p_{s}$ is both divergence-free and curl-free.
@@ 2.4 Vector Identities
A proof of the following theorem can be found in a standard text on differentiable manifolds such as [27].
Theorem 3 (Gauss-Green).
@@
Apply Theorem 3 to $uv$ and we have
Lemma 2 (Integration-by-parts).
@@
Lemma 3 (Green’s formula).
@@ 2.5 B-stable and algebraically stable RK methods
A function $\mathbf{f}:\mathbb{R}^{{\scriptsize\textup{D}}}\times[0,+\infty)\to\mathbb{R}^{{\scriptsize\textup{D}}}$ satisfies a one-sided Lipschitz condition if
@@
A contractive ODE system is dissipative: for any solutions $\mathbf{u}(t)$ and $\mathbf{v}(t)$, the norm $\|\mathbf{u}(t)-\mathbf{v}(t)\|$ decreases monotonically as $t$ increases. In other words, different solution trajectories of a contractive ODE system never depart from each other, and hence small perturbations remain small. For numerical methods, this leads to
Definition 1 (B-stability [28]).
A one-step method is B-stable if, for any contractive ODE system, every pair of its numerical solutions $\mathbf{U}^{n}$ and $\mathbf{V}^{n}$ satisfy
@@
An algebraically stable RK method is B-stable and thus A-stable [29].
@@ 3 The GePUP-E formulation
Boundary conditions of the evolutionary variable $\mathbf{w}$ play a crucial role in establishing a stable numerical scheme. As elaborated in Subsection 1.2.4, the GePUP formulation (21) leads to the heat equation (22). However, neither the homogeneous Dirichlet nor the homogeneous Neumann condition is explicitly imposed on $\nabla\cdot\mathbf{w}$ in (21), thus the dacay of $\nabla\cdot\mathbf{w}$ is not guaranteed. To fix this glitch, we draw inspiration from the excellent work of Rosales, Shirokoff, and their colleagues [22, 23] to adapt “electric” boundary conditions into GePUP, proposing
Definition 3.
The GePUP-E formulation of the INSE on no-slip domains is
$\frac{\partial\mathbf{w}}{\partial t}$ $=\mathbf{g}-\mathbf{u}\cdot\nabla\mathbf{u}-\nabla q+\nu\Delta\mathbf{w}$ in $\text{in }\Omega,$ (39a)
$\mathbf{w}\cdot\bm{\tau}$ $=0,\ \ \nabla\cdot\mathbf{w}=0$ on $\text{on }\partial\Omega,$ (39b)
$\mathbf{u}$ $={\mathscr{P}}\mathbf{w}$ in $\text{in }\Omega,$ (39c)
$\mathbf{u}\cdot\mathbf{n}$ $=0$ on $\text{on }\partial\Omega,$ (39d)
$\Delta q$ $=\nabla\cdot(\mathbf{g}-\mathbf{u}\cdot\nabla\mathbf{u})$ in $\text{in }\Omega,$ (39e)
$\mathbf{n}\cdot\nabla q$ $=\mathbf{n}\cdot(\mathbf{g}-\mathbf{u}\cdot\nabla\mathbf{u}+\nu\Delta\mathbf{w})+\lambda\mathbf{n}\cdot\mathbf{w}$ on $\text{on }\partial\Omega,$ (39f)
@@
(39b) is different from (21b). First, the boundary condition $\mathbf{n}\cdot\mathbf{w}=0$ in (21b) is removed and the term $\lambda\mathbf{n}\cdot\mathbf{w}$ is added in (39f) so that any nonzero $\mathbf{n}\cdot\mathbf{w}$ decays exponentially towards zero; see Lemma 5. Second, the boundary condition $\nabla\cdot\mathbf{w}=0$ is added in (39b) to set up an exponential decay of $\nabla\cdot\mathbf{w}$; see Theorem 4. Third, the boundary condition $\mathbf{w}\cdot\bm{\tau}=0$ is added in (39b) to close the vector diffusion equation governing the evolution of $\mathbf{w}$. Lastly, the boundary condition $\mathbf{u}\cdot\bm{\tau}=0$ in (21b) is removed because, as a perturbed version of $\mathbf{u}$, the non-solenoidal velocity $\mathbf{w}$ is $\mathbf{u}$ when the initial condition (40) is imposed; see Lemma 6. Another initial condition weaker than (40) is
@@
Compared to the formulation proposed in [22, 23], the GePUP-E formulation (39) facilitates the design and analysis of numerical schemes that treat the nonlinear convection term and the pressure gradient term explicitly; see Definition 6.
Lemma 4.
The Neumann BVP (39c,d) admits a unique solution of $\mathbf{u}$.
Proof.
(39c,d) and the definition $\mathbf{w}=\mathbf{u}-\nabla\phi$ yield
@@
Then the proof is completed by Theorem 1 and the divergence theorem. ∎

### v2

@@ 1.1.4 Generic projection and unconstrained PPE (GePUP)
To accommodate the fact that the discrete velocity might not be divergence-free, we switch the evolutionary variable to a non-solenoidal velocity $\mathbf{w}={\mathcal{P}}\mathbf{u}$ instead of the divergence-free velocity $\mathbf{u}$ in the UPPE formulation (13). More precisely, the evolutionary variable $\mathbf{u}$ in the time-derivative term $\frac{\partial\mathbf{u}}{\partial t}$ is perturbed to a non-solenoidal velocity $\mathbf{w}:=\mathbf{u}-\nabla\phi$ where $\phi$ is some scalar function; meanwhile in the diffusion term we change $\mathbf{u}$ to $\mathbf{w}$ to set up a mechanism that drives the divergence towards zero. Then, there is no need to worry about the influence of $\nabla\cdot\mathbf{w}\neq 0$ on numerical stability because the evolution of $\mathbf{w}$ is not subject to the divergence-free constraint. These ideas lead to the GePUP formulation zhang16:_GePUP :
$\frac{\partial\mathbf{w}}{\partial t}$ $=\mathbf{g}-\mathbf{u}\cdot\nabla\mathbf{u}-\nabla q+\nu\Delta\mathbf{w}$ $\text{in }\Omega,$ (15a)
$\mathbf{w}$ $=\mathbf{0},\ \ \mathbf{u}\cdot\bm{\tau}=0,\ \$$\text{on }\partial\Omega,$ (15b)
$\mathbf{u}$ $={\mathscr{P}}\mathbf{w}$ $\text{in }\Omega,$ (15c)
$\mathbf{u}\cdot\mathbf{n}$ $=0$ $\text{on }\partial\Omega,$ (15d)
$\Delta q$ $=\nabla\cdot(\mathbf{g}-\mathbf{u}\cdot\nabla\mathbf{u})$ $\text{in }\Omega,$ (15e)
$\mathbf{n}\cdot\nabla q$ $=\mathbf{n}\cdot(\mathbf{g}+\nu\Delta\mathbf{u}-\nu\nabla\nabla\cdot\mathbf{u})$ $\text{on }\partial\Omega,$ (15f)
where (15e) and the divergence of (15) yield $\frac{\partial\left(\nabla\cdot\mathbf{w}\right)}{\partial t}=\nu\Delta\left(\nabla\cdot\mathbf{w}\right)$. Then either $\nabla\cdot\mathbf{w}=0$ or $\mathbf{n}\cdot\nabla\nabla\cdot\mathbf{w}=0$ on $\partial\Omega$ drives $\nabla\cdot\mathbf{w}$ towards zero.
@@ 1.2 The contribution of this work
We couple GePUP with electric boundary conditions shirokoff11:_navier_stokes ; rosales21:_high_poiss_navier_stokes and a scalar auxiliary variable (SAV) shen18:SAV ; shen19:SAV_Review to propose GePUP-E and GePUP-ES, variants of GePUP that enforce the solenoidal conditions, preserve energy stability, decouple time integration from spatial discretization, and lead to versatile algorithms that are fourth-order accurate both in time and in space. The letter ‘E’ in the acronyms stands for the electric boundary conditions while the letter ‘S’ for the SAV approach; altogether ‘ES’ also stands for energy stability.
GePUP-E and GePUP-ES answer all the challenges in (FMC-1,2,3,4).
  1. (A)
    We reformulate the INSE of two variables into the GePUP-E formulation (32) of three variables, embedding the solution manifold $\mathcal{M}(\mathbf{u},p)$ of the INSE in the higher-dimensional solution manifold $\mathcal{N}(\mathbf{w},\mathbf{u},q)$ of GePUP-E. More importantly, we equip this embedding with a divergence-decaying mechanism that drives any deviating solution in $\mathcal{N}(\mathbf{w},\mathbf{u},q)$ back to $\mathcal{M}(\mathbf{u},p)$. We prove the equivalence of INSE and GePUP-E, the convergence of the non-solenoidal velocity $\mathbf{w}$ to the divergence-free velocity $\mathbf{u}$, the exponential decay of the divergence $\nabla\cdot\mathbf{w}$, and the monotonic decrease of the kinetic energy. As such, GePUP-E resolves the difficulties in (TMR-1,2).
  2. (B)
    By coupling GePUP-E to SAV shen18:SAV ; shen19:SAV_Review , we propose the GePUP-ES formulation in (47), design a family of semi-discrete GePUP-ES algorithms, and prove their energy stability in Theorem 5.4.
  3. (C)
    Based on (B), we further propose a family of fully discrete INSE solvers, named GePUP-ES-SDIRK, to answer all challenges in (FMC-1,2,3,4).
(A), (B), and (C) are elaborated in Sections 3, 4, and 5, respectively. In Section 2, we introduce notation to make this paper somewhat self-contained. We test GePUP-ES-SDIRK in Section 6 and draw conclusions in Section 7.
@@ 2 Preliminaries
Throughout this paper, we denote by $\left\langle\cdot,\cdot\right\rangle$ the $L^{2}$ inner product of vector- (or scalar-) valued functions $\mathbf{u}$ and $\mathbf{v}$ over $\Omega$, $\left\langle\mathbf{u},\mathbf{v}\right\rangle:=\int_{\Omega}\mathbf{u}\cdot\mathbf{v}\,\mathrm{d}V$, and by $\|\cdot\|$ the induced $L^{2}$ norm $\|\mathbf{u}\|:=\sqrt{\left\langle\mathbf{u},\mathbf{u}\right\rangle}$.
We start with a well-known result on boundary value problems (BVPs).
Theorem 2.1 (Solvability of BVPs with pure Neumann conditions)
@@
Proof
See (taylor11:_partial_differ_equat_i, , page 409). ∎
2.1 The Leray-Helmholtz projection $\mathscr{P}$}
Theorem 2.2 (Helmholtz decomposition)
@@ 2.1 The Leray-Helmholtz projection $\mathscr{P}$}
since Theorem 2.1 uniquely determines $\nabla\phi$ with $\oint_{\partial\Omega}\mathbf{v}\cdot\mathbf{n}=0$. ∎
@@
The proof of Theorem 2.2 implies the constructive form ${\mathscr{P}}={\mathcal{I}}-\nabla(\Delta_{n})^{-1}\nabla\cdot$, where $(\Delta_{n})^{-1}$ denotes solving (18). For a ${\mathcal{C}}^{1}$ vector field $\mathbf{v}^{*}$ and a ${\mathcal{C}}^{1}$ scalar field $\phi$, we have
@@ 2.2 The Laplace-Leray commutator $\Delta{\mathscr{P}}-{\mathscr{P}}\Delta$
On periodic domains, $\Delta$ and $\mathscr{P}$} commute. However, one main difficulty for no-slip domains is the fact that $\Delta{\mathscr{P}}-{\mathscr{P}}\Delta\neq\mathbf{0}$. In this subsection we rephrase several results in liu07:_stabil_conver_effic_navier_stokes .
Lemma 2.3
@@
satisfies $\nabla\cdot{\mathcal{B}}=0$, $\Delta{\mathscr{P}}={\mathcal{B}}$, and in three dimensions $\mathcal{B}$$\nabla\times\nabla\times$.
Proof
@@
Then Lemma 2.3 and (23) give
Corollary 2.4
@@
$\Delta,{\mathscr{P}}]:=\Delta{\mathscr{P}}-{\mathscr{P}}\Delta=({\mathcal{I}}-{\mathscr{P}})\Delta{\mathscr{P}}=({\mathcal{I}}-{\mathscr{P}}){\mathcal{B}}=-({\mathcal{I}}-{\mathscr{P}})(\nabla\times\nabla\times),$ (24)
@@ 2.3 The Stokes pressure
By (24), the action of the Laplace-Leray commutator on any vector field $\mathbf{v}^{*}$ results in the gradient of some scalar field. In the case of $\mathbf{v}^{*}$ being the velocity $\mathbf{u}$ in the INSE, the scalar is known as the Stokes pressure liu07:_stabil_conver_effic_navier_stokes :
@@
Then (20) and $\nabla\cdot{\mathcal{B}}=0$ in Lemma 2.3 yield $\Delta p_{s}=0$, i.e., the Stokes pressure is harmonic. Interestingly, the vector field $\nabla p_{s}$ is both divergence-free and curl-free.
@@ 2.4 Vector identities
See a standard text on differentiable manifolds such as munkres97:_analy_manif for a proof of
Theorem 2.5 (Gauss-Green)
@@
Apply Theorem 2.5 to $uv$ and we have
Lemma 2.6 (Integration-by-parts)
@@
Lemma 2.7 (Green’s formula)
@@ 2.5 B-stable and algebraically stable RK methods
A function $\mathbf{f}:\mathbb{R}^{{\scriptsize\textup{D}}}\times[0,+\infty)\to\mathbb{R}^{{\scriptsize\textup{D}}}$ is one-sided Lipschitz continuous if
@@
A contractive ODE system is dissipative: for any solutions $\mathbf{u}(t)$ and $\mathbf{v}(t)$, the norm $\|\mathbf{u}(t)-\mathbf{v}(t)\|$ decreases monotonically as $t$ increases. In other words, different solution trajectories of a contractive ODE system never depart from each other, and hence small perturbations remain small. This leads to
Definition 2.8 (B-stability butcher75:_runge_kutta )
A one-step method is B-stable if, for any contractive ODE system, each pair of numerical solutions $\mathbf{U}^{n}$ and $\mathbf{V}^{n}$ satisfy
@@
An algebraically stable RK method is B-stable and thus A-stable hairer96:_solvin_ordin_differ_equat_ii .
@@ 3 The GePUP-E formulation
Boundary conditions of the evolutionary variable $\mathbf{w}$ play a crucial role in establishing a stable numerical scheme. As mentioned in Subsection 1.1.4, the GePUP formulation (15) leads to the heat equation $\frac{\partial\left(\nabla\cdot\mathbf{w}\right)}{\partial t}=\nu\Delta\left(\nabla\cdot\mathbf{w}\right)$. However, neither the homogeneous Dirichlet nor the homogeneous Neumann condition is explicitly imposed on $\nabla\cdot\mathbf{w}$ in (15), thus it is difficult to prove the dacay of $\nabla\cdot\mathbf{w}$. To fix this glitch, we draw inspiration from the excellent work of Rosales, Shirokoff, and their colleagues shirokoff11:_navier_stokes ; rosales21:_high_poiss_navier_stokes to adapt “electric” boundary conditions into GePUP, proposing
Definition 3.1
The GePUP-E formulation of INSE on no-slip domains is
$\frac{\partial\mathbf{w}}{\partial t}$ $=\mathbf{g}-\mathbf{u}\cdot\nabla\mathbf{u}-\nabla q+\nu\Delta\mathbf{w}$ $\text{in }\Omega,$ (32a)
$\mathbf{w}\cdot\bm{\tau}$ $=0,\ \ \nabla\cdot\mathbf{w}=0\ \$$\text{on }\partial\Omega,$ (32b)
$\mathbf{u}$ $={\mathscr{P}}\mathbf{w}$ $\text{in }\Omega,$ (32c)
$\mathbf{u}\cdot\mathbf{n}$ $=0$ $\text{on }\partial\Omega,$ (32d)
$\Delta q$ $=\nabla\cdot(\mathbf{g}-\mathbf{u}\cdot\nabla\mathbf{u})$ $\text{in }\Omega,$ (32e)
$\mathbf{n}\cdot\nabla q$ $=\mathbf{n}\cdot(\mathbf{g}-\mathbf{u}\cdot\nabla\mathbf{u}+\nu\Delta\mathbf{w})+\lambda\mathbf{n}\cdot\mathbf{w}$ $\text{on }\partial\Omega,$ (32f)
@@
(32b) is different from (15b). First, the boundary condition $\mathbf{n}\cdot\mathbf{w}=0$ in (15b) is removed and the term $\lambda\mathbf{n}\cdot\mathbf{w}$ is added in (32f) so that any nonzero $\mathbf{n}\cdot\mathbf{w}$ decays exponentially towards zero; see Lemma 3.3. Second, the boundary condition $\nabla\cdot\mathbf{w}=0$ is added in (32b) to set up an exponential decay of $\nabla\cdot\mathbf{w}$; see Theorem 3.4. Third, the boundary condition $\mathbf{w}\cdot\bm{\tau}=0$ is added in (32b) to close the vector diffusion equation governing the evolution of $\mathbf{w}$. Lastly, the boundary condition $\mathbf{u}\cdot\bm{\tau}=0$ in (15b) is removed because, as a perturbed version of $\mathbf{u}$, the non-solenoidal velocity $\mathbf{w}$ is $\mathbf{u}$ when the initial condition (33) is imposed; see Lemma 3.5. Another initial condition weaker than (33) is
@@
Compared to the formulation in shirokoff11:_navier_stokes ; rosales21:_high_poiss_navier_stokes , the GePUP-E formulation (32) facilitates the design and analysis of numerical schemes that treat the nonlinear convection term and the pressure gradient term explicitly; see Definition 5.1.
Lemma 3.2
The Neumann BVP (32c,d) admits a unique solution of $\mathbf{u}$.
Proof
(32c,d) and the definition $\mathbf{w}=\mathbf{u}-\nabla\phi$ yield
@@
Then the proof is completed by Theorem 2.1 and the divergence theorem. ∎